OpenStudy (aaronandyson):

A footpath of uniform width runs around the inside of a rectangular field 32 m long and 24 m wide.If the path occupies 208 m^2 ,find the width of the path

OpenStudy (er.mohd.amir):

let footpath length be x then length = 2x+24 and width =2x+32 area = (2x+24)*(2x+32)=208 this become quadratic in x solve for x

OpenStudy (aaronandyson):

???

OpenStudy (aaronandyson):

@perl @amistre64

OpenStudy (amistre64):

well, how would you propose we measure the area of the foot path to begin with?

OpenStudy (aaronandyson):

Can you be more specific ?

OpenStudy (amistre64):

me? no, that was as specific as i could get ..

OpenStudy (shamim):

Cam u draw figure? @Aron

OpenStudy (aaronandyson):

@amistre64 I did not really understand your statement/post.

OpenStudy (amistre64):

i asked for you thoughts about how you would attempt the solution. you must have some ideas to work with

OpenStudy (amistre64):

Er has an idea, but im not sure if its correctly expressed.

OpenStudy (aaronandyson):

Unable to understand :/ his solution

OpenStudy (shamim):

32x*2+(24-2x)x*2=208 Is it make any sence

OpenStudy (amistre64):

start with drawing a picture

OpenStudy (aaronandyson):